# VoCoT: Unleashing Visually Grounded Multi-Step Reasoning in Large Multi-Modal Models

### v2

@@ 5.2 Multi-Step Reasoning within Large Pre-trained Models
To address complex multi-modal problems, Shikra [6] initially explores the potential of applying CoT to specific tasks with LMMs. SoM [54] and Scaffolding [20] incorporate additional structural information into images, such as segmentation maps and dot grids. These enhancements help the model utilize these anchors for reasoning. However, these approaches are specifically designed for GPT-4V, which has strong visual perception capabilities. V-Star [53] designs a reasoning logic for visual search in high-resolution images, which can be considered a simple two-step CoT. Overall, CoT has not been comprehensively explored in LMMs and there lack appropriate reasoning formats that could be generalized to various scenarios and tasks.
@@ 6 Conclusion
In this paper, we introduce VoCoT, a visually-grounded and object-centric Chain of Thoughts (CoT) format, designed to assist large multimodal models (LMMs) in multi-step reasoning. Furthermore, we propose a pipeline to construct VoCoT-formatted data from existing resources, leading to the creation of VoCoT-Instruct-80K. Based on this data, we develop VolCano, a model designed to perform multi-step reasoning effectively and efficiently using the VoCoT format. Comprehensive experiments demonstrate the effectiveness of our approach.
@@ Appendix A Supplementary Details
A.1 Data Construction Details
Construction Methods
As for the Type 1 data construction, we utilize a rule-based conversion method, the mapping rules used are listed in Table 11. In terms of Type 2 and Type 3 data, the prompts for GPT-4V are respectively shown in Table 12 and  13. We use in-context learning to let GPT-4V generate thought with multi-step reasoning path in VoCoT format.
Quality Control
For Type 1 data generated by rule based mapping, the quality is controled by the original source, namely GQA. We do not perform additional quality control. For Type 2 and Type 3 data generated by GPT-4V, we manually check 200 samples, we find if GPT-4V think the information we provide is not enough to generate correct reasoning path, model response will contains error messages. All error message have patterns in common which may contains following phrase: “From the object information provided", “provided object information", “From the bounding boxes provided", “provided bounding boxes", “provided object bounding boxes", “object bounding boxes provided", “object bounding box information provided", “bounding box information provided", “object information provided", “given object information", “the bounding boxes provided", “From the object bounding box information", “provided object bounding box information". We remove the sample containing these patterns, leaving 8k samples after the filtering.
A.2 Training Data Details
Table 7: Data Mixture in 3 Stages of VolCano.
Stages Data Type Source Size
Stage 1 Image-Caption LLaVA-Pretrain [31; 29] 558k
Stage 2 Image-Caption ALLaVA-Caption [4] 695k
Grounded Image Caption GRIT [38] 756k
Flickr30k-Entities [39] 148k
Multimodal Document MMC4 [63] 890k
Stage 3 Visual Instruction LLaVA [29] 612k
Referring Expression Shikra-RD [6] 6k
RefCOCO [18] 42k
RefCOCO+ [18] 42k
RefCOCOg [18] 42k
g-RefCOCO [13] 79k
VoCoT This Work 80k
We present our data mixture details in Table 7. In Stage 1, we use LLaVA-pretrain [29] dataset for projector alignment which contains 558k image caption pairs. In Stage 2, to better adapt to multi-modal interleaved sequences and visually grounded object representations with Refbind, we mix three types of data: (1) multimodal documents, (2) grounded image captions, and (3) high-quality image captions. Multimodal documents data is sample from MMC4 [63] by choosing which average similarity score between image and sentence before each image is larger than 0.3. Each multimodal document have multiple image, we remove samples with more than 6 images. Grounded image captions is from GRIT [38] and Flickr30K Entities [39]. For GRIT, we filter samples with clip score is larger than 0.35. We also use high quality image caption from ALLaVA [4] which is generated by GPT-4V and provide details description. In Stage 3, we remove samples from LLaVA-Instruct that meet two criteria: sourced from RefCOCO, and sourced from VG where the object sub-image size is less 50, the reason is that we find extremely small regions in VG are probably with low quality.
A.3 Model & Training Details
We supplement the hyper-parameters in each stage in Table 8.
Table 8: The detailed training hyper-parameters of VolCano
Configuration Alignetent Multi-modal Interleaved Instruction Tuning
@@ A.3 Model & Training Details
Peak learning rate of LLM 5e-5 1e-5 1e-5
@@
Following Kosmos2 [38], we introduce a special token “<grounding>” at the beginning of the sequence to control whether to require the model to produce visually grounded description.
Also notice that VolCano is able to perform both single-step and multi-step reasoning. Following [19], we introduce a prompt “Answer the question and include the reasoning proess. Locate key objects and provide bounding boxes in your thoughts.” as a trigger to tell the model whether to use VoCoT or not during the generation process.
@@ A.4 Evaluation Details
Category General VQA Spatial Reasoning Hallucination
Datasets GQA MMBench SEED VSR EmbSpat. POPE AMBER
Data Amount 12578 4329 14233 1222 3625 3000 1004
@@
Category Compositional Tasks Referring Expression
Datasets V-Star Wino CLEVR RefCOCOg CLEVR-Ref
Split - test val val (umd) test (umd) -
Data Amount 238 800 6000 4896 9602 2000
@@ A.4.1 Benchmark Details
In GQA, we utilize the “testdev_balanced” split following [31]. As for MMBench, we adopt the “DEV” split for the evaluation efficiency. In terms of SEED, we only consider the subset that the visual inputs are images. For GQA, we append a prompt “Please answer in a word or short phrase” to require models produce concise outputs.
@@
For POPE, we consider the adversarial subset since it is the most challenging split. In AMBER, we leverage the generative task which asks the model to describe the image. All prompts are directly adopted from the original datasets.
Benchmarks for Compositional Tasks
For CLEVR, we utilize the val split. Because the original CLEVR validation set is too large, we categorize the data the into six types based on the question type: count, yes/no, shape, material, size, and color. We sample 1k questions from each category as test samples and construct a multiple-choice candidate set based on the feasible answers in the dataset. We will also open-source this subset. For Wino, we utilize the test set and consider it as a caption selection multiple-choice question, the prompt is designed as “Please describe the image.”. As for V-Star, we directly use the V-Star benchmark.
Referring Expression Benchmarks
For RefCOCOg, we utilize the val and test splits which is provided by umd. As for CLEVR-Ref, we use the provided set for evaluation. We design a prompt as "Can you locate phrase in the image?" where “{phrase}” is the target query.
Please see Table 10 and Table 9 for the dataset splits and scales.
@@ A.4.2 Evaluation Methods
For multiple-choice questions, we utilize the likelihood-based evaluation method, which is also known as the perplexity-based method. These methods are widely adopted in evaluating LMMs [21; 22; 9; 25]. The key idea is to select the option with the highest generated likelihood, please refer to these papers for the detail. If VoCoT is utilized, the likelihood is computed based on the question, image, and the genrated reasoning path.
@@
For GQA, we use the evaluation script provided by LLaVA [29] for a fair comparison. As for VSR and POPE, we require the model to answer in yes and no, enabling us to evaluate the correctness with exact match. With respoect to AMBER, we use the official evaluation method to assess the hallucinations in the generated descriptions.
@@
We first extract the predicted boxes from the outputs based on rules, then IoU between the ground truth box and predicted box is calculated. If the IoU is larger than 0.5, it is considered as a correct prediction [18].
@@
Appendix B Supplementary Analysis and Discussion
B.1 Case Study

### v3

@@ 5.2 Multi-Step Reasoning
To address complex multi-modal problems, Shikra Chen et al. (2023b) initially explores the potential of applying CoT to specific tasks with LMMs. SoM Yang et al. (2023a) and Scaffolding Lei et al. (2024) respectively incorporate segmentation maps and dot grids in images to assist LMMs in reasoning, but such information can only be utilized by proprietary models like GPT-4V. The most related work is VisCoT Shao et al. (2024), which designs a two-step CoT: first searching for a relevant region and then answering based on the additional region information. This method is effective but cannot model complex multi-step reasoning. Overall, CoT has not been comprehensively explored in LMMs and there lack appropriate reasoning formats that could be generalized to various scenarios and tasks.
@@ 6 Conclusion
In this paper, we introduce VoCoT, a visually-grounded and object-centric chain of thoughts format to assist LMMs in multi-step reasoning. We also curate a VoCoT-formatted dataset from existing resources to train LMMs to learn reasoning with VoCoT. Building on this, we develop VolCano, a model capable of multi-step reasoning using the VoCoT format. Comprehensive experimental results demonstrate the effectiveness of our approach.
Limitations
Our work, as an early exploration of CoT techniques in large multi-modal models, is limited in the following aspects. (1) Currently, VoCoT is designed for single-image context and not applicable to multi-image inputs like videos and image sequences. Additional special tokens or marks can be introduced to extend VoCoT to a two-step grounding for multiple images. Each object is first grounded to a specific image and then localized to a region within that image. We will explore such mechanisms in our future work. (2) The construction of VoCoT-formatted dataset is limited by the cost of calling proprietary models and can not effectively scale up. In future work, we will explore methods to reduce the cost of data construction, including using smaller or open-source models, collecting and converting more finely annotated data (such as DocVQA) in a manner similar to GQA, and simulating and generating data based on specific needs, similar to CLEVR. (3) The presented VolCano model is currently limited with respect to 7B-sized models due to the lack of computational resources. As implied by the experimental results in Section 4.5, we hope to demonstrate the potential of applying VoCoT to larger and stronger backbones as explored in textual CoT techniques.
Ethical Statement
The presented VoCoT-Instruct-80K dataset is sourced from open-source datasets including GQA Hudson and Manning (2019), LLaVA-Instruct Liu et al. (2024b), and LVIS Gupta et al. (2019). We carefully follow the license to use these datasets and ensure that they are applicable for research purposes. The original datasets have been widely adopted by relevant researchers and ensure no risk of privacy leakage or harmful information. Furthermore, during the data collection and construction, we perform balanced sampling based on the distribution of object categories to alleviate distribution bias. As mentioned in Appendix A.1, we also conduct human-in-the-loop quality control to ensure the final dataset has correct information without ethical issues. Please refer to Appendix B.9 for the detailed discussion. Currently, the presented models and dataset focus on English, we hope to expand to other languages in the future. Our work and artifacts are designed with the principle of universality and fairness, without any preference for specific demographic groups.
Acknowledgment
The work is supported by National Key R&D Program of China (Grant Nos. 2023YFF1204800) and National Natural Science Foundation of China (Grant Nos. 62176058). The project’s computational resources are supported by CFFF platform of Fudan University.
@@ Appendix A Supplementary Details
Configuration Alignment Multi-modal Interleaved Instruction Tuning
@@
Peak learning rate of LLM 1e-3 1e-5 1e-5
@@
Table 7: The detailed training hyper-parameters of VolCano. Except for the backbones for initialization, VolCanoQ2 follows the same hyper-parameters.
A.1 Data Construction Details
Construction Methods
As for the Type 1 data construction, we utilize a rule-based conversion method, the mapping rules used are listed in Table 16. In terms of Type 2 and Type 3 data, the prompts for GPT-4V are respectively shown in Table 17 and  18. We use in-context learning to let GPT-4V generate thought with multi-step reasoning path in VoCoT format.
Quality Control
For Type 1 data generated by rule based mapping, the quality is controled by the original source, namely GQA. We do not perform additional quality control.
For Type 2 and Type 3 data generated by GPT-4V, we first perform balanced sampling for images to achieve a balanced distribution of objects included. Secondly, we manually sample and check 200 data samples and find that initially constructed data suffers issues like uneven question types and incorrect object information. The first issue can be addressed by including well-designed in-context samples. The second issue is mainly caused by the potential incomplete and incorrect labels in LVIS. We find if GPT-4V think the information we provide is not enough to generate correct reasoning path, the response will contains error messages. All error message have patterns in common which may contains the following phrases: “From the object information provided", “provided object information", “From the bounding boxes provided", and so on. We remove the samples containing these patterns, leaving 8k samples after the filtering. Ultimately, by manually checking the constructed dataset, we do not observe any bias issues and achieve a 98% pass rate on the presented dataset.
A.2 Training Data Details
Stages Data Type Source Size
Stage 1 Image-Caption LLaVA-Pretrain Liu et al. (2024b, 2023b) 558k
Stage 2 Image-Caption ALLaVA-Caption Chen et al. (2024) 695k
Grounded Image Caption GRIT Peng et al. (2023) 756k
Flickr30k-Entities Plummer et al. (2015) 148k
Multimodal Document MMC4 Zhu et al. (2024) 890k
Stage 3 Visual Instruction LLaVA Liu et al. (2023b) 612k
Referring Expression Shikra-RD Chen et al. (2023b) 6k
RefCOCO Kazemzadeh et al. (2014) 42k
RefCOCO+ Kazemzadeh et al. (2014) 42k
RefCOCOg Kazemzadeh et al. (2014) 42k
g-RefCOCO He et al. (2023) 79k
VoCoT This Work 80k
Table 8: The data mixture used in the three training stages of VolCano.
We present our data mixture details in Table 8. In Stage 1, we use LLaVA-pretrain Liu et al. (2023b) dataset for projector alignment which contains 558k image caption pairs. In Stage 2, to better adapt to multi-modal interleaved sequences and visually grounded object representations with Refbind, we mix three types of data: (1) multimodal documents, (2) grounded image captions, and (3) high-quality image captions. Multimodal documents data is sample from MMC4 Zhu et al. (2024) by choosing which average similarity score between image and sentence before each image is larger than 0.3. Each multimodal document have multiple image, we remove samples with more than 6 images. Grounded image captions is from GRIT Peng et al. (2023) and Flickr30K Entities Plummer et al. (2015). For GRIT, we filter samples with clip score is larger than 0.35. We also use high quality image caption from ALLaVA Chen et al. (2024) which is generated by GPT-4V and provide details description. In Stage 3, we remove samples from LLaVA-Instruct that meet two criteria: sourced from RefCOCO, and sourced from VG where the object sub-image size is less 50, the reason is that we find extremely small regions in VG are probably with low quality.
A.3 Model & Training Details
The hyper-parameters in each stage are in Table 7. The learning rate setup mainly follows that of LLaVA Liu et al. (2024b). In stage 1, a large learning rate is used to update the connection module, aiming to quickly align the cross-modal representations. In the latter stages, a small learning rare is adopted to carefully fine-tune the backbone. Following Kosmos2 Peng et al. (2023), we introduce a special token “<grounding>” at the beginning of the sequence to control whether to require VolCano to produce visually grounded description.
Also notice that VolCano is able to perform both single-step and multi-step reasoning. Following Kojima et al. (2022), we introduce a prompt “Answer the question and include the reasoning proess. Locate key objects and provide bounding boxes in your thoughts.” as a trigger to tell the model whether to use VoCoT or not during the generation process.
@@ A.4 Evaluation Details
Category General VQA Spatial Reasoning Hallucination
Dataset GQA MMBench SEED VSR EmbSpat. POPE AMBER
Size 12578 4329 14233 1222 3625 3000 1004
@@
Category Composite Tasks Referring Expression
Dataset V-Star Wino CLEVR CLEVR-Ref
Split - test val -
Size 238 800 6000 2000
@@ A.4.1 Benchmark Details
In GQA, we utilize the “testdev_balanced” split following Liu et al. (2024b). As for MMBench, we adopt the “DEV” split for the evaluation efficiency. In terms of SEED, we only consider the subset that the visual inputs are images. For GQA, we append a prompt “Please answer in a word or short phrase” to require models produce concise outputs.
@@
For POPE, we consider the adversarial subset since it is the most challenging split. In AMBER, we leverage the generative task which asks the model to describe the image. All prompts are adopted from the original datasets with a yes-or-no instruction for POPE.
Benchmarks for Composite Tasks
For CLEVR, we utilize the val split. Because the original CLEVR validation set is too large, we categorize the data the into six types based on the question type: count, yes/no, shape, material, size, and color. We sample 1k questions from each category as test samples and construct a multiple-choice candidate set based on the feasible answers in the dataset. We will also open-source this subset. For Winoground, we utilize the test set and consider it as a caption selection multiple-choice question, the prompt is designed as “Please describe the image.”. As for V-Star, we directly use the V-Star benchmark. Regarding CLEVR-Ref, which is a referring expression task with relatively complex queries, we use the provided set for evaluation. We design a prompt as "Can you locate {phrase} in the image?" where “{phrase}” is the target query.
Please see Table 10 and Table 9 for the splits and scales of benchmarks used in this paper.
@@ A.4.2 Evaluation Methods
For multiple-choice questions, we utilize the likelihood-based evaluation method, which is also known as the perplexity-based method. These methods are widely adopted in evaluating LMMs Li et al. (2023a, b); Dai et al. (2023); Li et al. (2023e). The key idea is to select the option with the highest generated likelihood, please refer to these papers for the detail. If VoCoT is utilized, the likelihood is computed based on the question, image, and the genrated reasoning path.
@@
For GQA, we use the evaluation script provided by LLaVA Liu et al. (2023b) for a fair comparison. As for VSR and POPE, we require the model to answer in yes and no, enabling us to evaluate the correctness with exact match. With respoect to AMBER, we use the official evaluation method to assess the hallucinations in the generated descriptions.
@@
We first extract the predicted boxes from the outputs based on rules, then calculate the IoU between the ground truth box and the predicted box. If the IoU is larger than 0.5, it is considered as a correct prediction following  Kazemzadeh et al. (2014).
@@
A.5 Introduction to Baseline Models
We compare VolCano to several existing SOTA open-source LMMs, including BLIP-2 Li et al. (2023b), InstructBLIP Dai et al. (2023), Shikra Chen et al. (2023b), mPLUG-Owl2 Ye et al. (2023), MiniGPT-v2 Chen et al. (2023a), Qwen-VL-Chat Bai et al. (2023), VILA Lin et al. (2024), LLaVA-1.5 Liu et al. (2023b), and the most related VisCOT Shao et al. (2024). These models are based on baseline LLM backbones released in 2023, including LLaMA Touvron et al. (2023a), LLaMA-2 Touvron et al. (2023b), Vicuna Chiang et al. (2023), Mistral Jiang et al. (2023), and Qwen Bai et al. (2023). For models based on recently proposed advanced backbones like LLaMA-3 Dubey et al. (2024) and Qwen2 Yang et al. (2024), we compare VolCanoQ2 with Bunny He et al. (2024) and VILA-1.5 Lin et al. (2024). The models listed before take a single image as input, which is consistent with VolCano and VolCanoQ2 for a fair comparison. Additionally, we include another series of SOTA LMMs that enhance input resolution by splitting a single image into multiple sub-images: LLaVA-1.6 Liu et al. (2024a), Deepseek-VL Lu et al. (2024), and Monkey Li et al. (2023f). As for GPT-4V, examples in Figure 1 and all results are obtained by calling openai API using the “GPT-4V” model between 2024/05/18 to 2024/05/30.
We only consider zero-shot performance in Table 2, except for GQA. If a model has been trained on a specific evaluation benchmark, we do not report the corresponding evaluation results. For example, in Table 2, the result of VisCOT on VSR is omitted because it uses the corresponding training data. For certain models, including Bunny and VILA, due to the lack of clear evaluation details, we re-evaluate their performance in the same setting to make a fair comparison with VolCano.
Appendix B Supplementary Results and Discussion
B.1 Comparison between VolCano and High-Resolution models
Model General VQA Spatial Reasoning Composite Tasks Hallucination
[…]
Models with single-image inputs
Models with multiple-image inputs
GPT-4V $2048^{2}$ 75.80 71.60 68.24 26.65 51.90 55.00 83.75 82.00 4.60
Table 11: Comparison between VolCano and resolution-enhanced models. The notations follows Table 2. LLaVA1.6-7Bm represents the Mistral-based LLaVA1.6-7B.
Model Scene-Text-Oriented Tasks Know. Reasoning Ins. Following Captioning
TextVQAN AI2D ChartQA DocVQA MMMU MathVista MMVet MMT COCO NoCaps
LLaVA1.5-7B 46.1 43.0 14.9 2.9 28.4 26.7 30.5 45.0 94.5 95.6
VolCano-SE 45.1 45.3 19.0 4.6 29.3 27.8 32.1 44.1 81.0 90.6
VolCano 48.9 45.6 19.5 10.6 33.3 27.9 32.9 45.4 100.4 103.5
Table 12: Peformance on additional benchmarks. TextVQAN represents the TextVQA benchmark without providing reference OCR like in LLaVA-1.5 Liu et al. (2023b). “Know.” and “Ins.” are respectively short for knowledge and instruction. The best performance for each dataset are bolded.
In Table 2 we compare models with single-image and relatively low-resolution inputs. Comparing VolCano with LMMs that enhance input resolution by introducing multiple-image inputs, we observe that these methods primarily improve the performance in general VQA and V-Star, as V-Star provides high-resolution input images Wu and Xie (2023). However, in tasks that require complex reasoning, the improvement brought by higher resolutions becomes less significant. VolCano either exceeds or perform comparably with these models in such tasks, indicating the superiority of introducing multi-step reasoning over enriching the input information in these scenarios.
Notice that the RefBind mechanism introduced in Section 3 can be directly extended to fit multiple split sub-images by mapping the predicted coordinates to patches from different sub-images. We leave exploring combining these two vertical research directions—enhancing input resolution and introducing multi-step reasoning–as future work.
Figure 7: Cases on additional benchmarks included in Appendix B.2.
B.2 Performance in Additional Benchmarks
The benchmarks presented in the main text primarily focus on object-centric scenarios, which align with the design of VoCoT. Recently, a line of research develops LMMs to handle scene-text-oriented scenarios including document understanding and chart information extraction. In this section, we explore whether VoCoT can adapt to other scenarios by conducting experiments on additional benchmarks: TextVQA Singh et al. (2019), AI2D Kembhavi et al. (2016), ChartQA Masry et al. (2022), and DocVQA Mathew et al. (2021) for scene-text-oriented tasks; MMMU Yue et al. (2023) and MathVista Lu et al. (2023) for reasoning based on knowledge and scene texts; MMVet Yu et al. (2023) and MMT Ying et al. (2024) for instruction-following; COCO caption Veit et al. (2016) and NoCaps Agrawal et al. (2019). For image captioning, we report the CIDEr Vedantam et al. (2015) metric following Li et al. (2023b) while accuracy is provided for other tasks. For efficiency, we only consider LLaVA-1.5, VolCano-SE, and VolCano for a fair and straightforward comparison to validate the effect of VoCoT.
Scene-Text-Oriented Benchmarks
As shown in Table 12, we can find: (1) Quantitatively, VoCoT can generalize to text and chart-centric scenarios and improve the performance, even though VoCoT-Instruct-80K does not include similar data. (2) Qualitatively, we present several samples in Figure 7. VoCoT can treat text blocks and chart areas as "objects". By locating and analyzing the corresponding regions, VoCoT easily generalizes to such tasks. (3) The cross-domain improvements are very exciting, and the potential of VoCoT in such tasks can be further unleashed by enhancing input resolution and constructing text-centric VoCoT data (with fine-grained annotations in TextVQA, AI2D…), which we leave as our future work.
Method Avg. # Tokens Generated Avg. # Visual Tokens Added Avg. Inference Time per Query
VolCano-SE 2.2 0 0.11s
VolCano 115.9 326 0.74s
Table 13: Statistics of single-step and multi-step inference in VSR.
Model MMBench SEED MMMU MMT MMVet
LLaVA1.5-7B 64.3 53.8 28.4 45.0 30.5
VolCano-SE 61.1 54.2 29.3 45.1 32.1
VolCano 63.3 57.5 31.2 45.2 32.5
Table 14: Performance of differnt models with single-step inference settings.
Knowledge Reasoning and Instruction Following
According to results in Table 12, we observe that: (1) Although MMMU and MathVista focus on reasoning involving commonsense, knowledge, and mathematical deduction, the results indicate that the generalization ability of VoCoT can still improve the performance. As shown in Figure 7, the framework of VoCoT to locate and analyze local objects and regions can generalize to such tasks. (2) Improvements brought by VoCoT in the absence of similar training data demonstrate its potential in such scenario, inspiring us to incorporate both visually grounded information and conceptual knowledge in VoCoT framework in our future work. (3) In general instruction-following datasets, MMT and MMVet, it is shown that VoCoT does not hurt the instruction-following ability and brings improvement, which is consistent with the findings in SEED and MMBench mentioned in our main paper. See Appenidx B.3 for further exploration.
Image Captioning
Results in Table 12 imply that VoCoT helps VolCano to perceive accurate information and produce visually grounded descriptions, improving the quality of generated captions.
Generally, the results validates that VoCoT can generalize across various tasks and demonstrates potential for further enhancement.
B.3 Do VoCoT Damage the Original Capabilities of LMMs?
Another concern about whether the proposed VoCoT framework brings negative effects to the original capabilities of a LMM. Noticing that the trained VolCano can perform both single-step and multi-step reasoning as mentioned in Appendix A.3 and our paper aims at unleashing VoCoT reasoning ability of LMMs without affecting the original abilities. To validate that, we evaluate VolCano with the same single-step inference strategy as other compared baselines on general benchmarks in Table 14. It can be seen that VolCano maintains the original ability for conventional single-step inference, performing close to the two baselines in such settings and even surpassing them in some tasks. This indicates that the VoCoT framework does not hurt the original and fundamental capabilities of LMMs.
B.4 Grounding Capabilities of VolCano
Besides the reasoning capability analyzed in Section 4.5, another key capability to ensure the effectiveness of VoCoT is the grounding. However, the commonly adopted RefCOCO Kazemzadeh et al. (2014) dataset is widely used in training LMMs and not applicable for zero-shot evaluation. Besides, another problem with RefCOCO is that the query is relatively simple. To address this, we consider CLEVR-Ref in the main part because complex queries are considered. As a step further, we are interested in the grounding capability of LMMs during the generation process, but it is difficult to evaluate under this setting. Therefore, we conduct a preliminary exploration.
Specifically, we evaluate the performance of models to produce grounded captions: requiring models to annotate objects with coordinates while describing the images. 100 images are sampled from the LVIS Gupta et al. (2019) validation set for evaluation. Pairwise evaluation is performed to compare the grounded contents generated by two models. Given the image, ground-truth object information, and responses from 2 models, the judge, GPT-4V, will score 2 responses from multiple perspectives (including both content accuracy and coordinates accuracy) and determine the winner.
Model LLaVA1.5 LLaVA1.5 VolCano
Visual Input Image Image + Object Info. Image
CLEVR Acc. 43.73 45.70 56.17
Table 15: Performance of models in CLEVR with different visual inputs. Acc. and Info. are short for Accuracy and Information, respectively.
As GPT-4V itself can not generate precise coordinates, we conduct a sanity check whether GPT-4V can evaluate the relevance between two bounding boxes described in texts. for ease of testing, we ask GPT-4V to evaluate responses in RefCOCOg that include single coordinates, and measure the Pearson correlation coefficient between "coordinate accuracy" judged by GPT-4V and the actual IoU. The coefficient is 0.932, indicating that GPT-4V can accurately judge the matching degree between coordinates represented in text. For ease of testing, we ask GPT-4V to evaluate responses in RefCOCOg that include single coordinates, and measure the Pearson correlation coefficient between "coordinate accuracy" judged by GPT-4V and the actual IoU. The coefficient is 0.932, indicating that GPT-4V can accurately judge the matching degree between coordinates represented in text.
Generally, we believe that the current setup can, to some extent, reflects the grounding abilities of models during generation. Ultimately, we compare VolCano with Qwen-VL-Chat and MiniGPTv2. The win rates of VolCano against Qwen-VL-Chat and MiniGPTv2 are 76.5 and 82.0, respectively, indicating that VolCano can perform better in simultaneously locating and describing visual contents.
B.5 Computational Efficiency
VoCoT leads to additional computational overheads compared with traditional single-step reasoning: (1) RefBind only introduces indexing operations without float-point calculations. Additional cost is caused by the visual tokens added to the sequences that will be processed. (2) Multi-step reasoning leads to additional computation by requiring more tokens to be generated. Since precise calculation of the computation cost is challenging, we provide empirical statistics in Table 13.
Notice that additional computation is inevitable in CoT methods. However, with Flash Attention Dao et al. (2022) and KV Cache methods used during generation, we found the increase in token quantities does not lead to excessive inference time. In the future, we will follow text CoT papers to explore efficient decoding methods which help improve the efficiency of VoCoT.
B.6 Can Other Open-Source LMMs Directly Utilize Object Information?
Besides the traditional single-step reasoning paradigm, we wonder whether open-source LMMs like LLaVA can utilize the provided ground truth object information to perform grounded reasoning and enhance the improvement. We conduct a experiment on CLEVR where gold object coordinates exist in the dataset. According to Table 15, it is observed that LLaVA benefits from the information but can not utilize it effectively. In contrast, VolCano can perform localization, analysis and reasoning on its own, showing clear superiority.
B.7 Potential Language Bias in Spatial Reasoning
As presented in 3, text-only CoT method performs the best in VSR, we attribute this phenomenon to two reasons: (i) scenarios in VSR are relatively simple, and (ii) the text-only models can better leverage the language bias in spatial relationships as a shortcut. Firstly, each type of object appears only once in an image, with a one-to-one correspondence between the text and the object. so coordinates are not required to resolve ambiguity.
Secondly, as noted in Kamath et al. (2023), spatial reasoning datasets exhibit some language biases (e.g., a television is more likely to be on a table rather than under it). We find that such biases are more likely to be exploited by text-only CoT models, while VoCoT-based VolCano relies more on analyzed visual information.
We conduct a experiment: replacing the images in VSR with completely black images and using the original queries to ask whether the corresponding spatial relationships exist in the image: (1) The VoCoT-based VolCano predicts "no" for 99% of the samples (in line with expectations). (2) The text-only CoT-based model predicted "yes" for 26% of the samples, achieving a 54.1% accuracy rate among these predictions (better than random choice for the binary questions). (3) This phenomenon demonstrates that the text-only CoT-based model is more prone to being influenced by language bias, which provides it with a shortcut and additional advantage in VSR.
B.8 Case Study